Circular No. 8924

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

CBAT@CFA.HARVARD.EDU (science)

URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 E2 = P/2003 KV_{2} (LINEAR) After the posting on the Minor Planet Center’s ‘NEOCP’ webpage of

an apparently asteroidal object discovered by the Catalina Sky Survey, S. Foglia (Novara Veveri, Italy) suggested an identification with comet P/2003 KV_{2} (cf. IAUC 8139), indicating a correction of ΔT = +0.8 day to the prediction by Nakano on MPC 56801. CCD astrometry:

2008 UT α_{2000} δ_{2000} Mag. Observer
Mar. 3.46904 15^{h}57^{m}40.^{s}93 +24^{o}18′15′′.2 18.7 Hill
3.47646 15 57 42.30 +24 18 14.6 19.0  _{”}
3.48384 15 57 43.52 +24 18 14.4 18.7  _{”}
3.49127 15 57 44.89 +24 18 14.4 19.0  _{”}
3.99893 15 59 12.01 +24 17 35.6 19.2 Süssenberger
4.00323 15 59 12.73 +24 17 34.9 19.5  _{”}
4.01054 15 59 13.98 +24 17 34.1 19.2  _{”}
4.04198 15 59 19.28 +24 17 32.4 20.0  _{”}
4.06552 15 59 23.34 +24 17 30.4  Birtwhistle
4.07005 15 59 24.09 +24 17 30.0  _{”}
4.07465 15 59 24.86 +24 17 30.1 19.5  _{”}
4.35803 16 00 13.55 +24 17 13.6 18.7 McGaha
4.36091 16 00 14.16 +24 17 12.7 19.2  _{”}
4.36378 16 00 14.54 +24 17 12.7 19.8  _{”}
4.48099 16 00 34.27 +24 17 05.5  Ries
4.48764 16 00 35.40 +24 17 04.9  _{”}
4.49428 16 00 36.48 +24 17 04.7  _{”}

R. E. Hill (Catalina). 0.68-m Schmidt telescope. U. Süssenberger (Bergen-Enkheim Observatory). 0.36-m reflector. P. Birtwhistle (Great Shefford, U.K.). 0.40-m reflector. J. E. McGaha (Sabino Canyon Observatory, Tucson). 0.36-m reflector. J. G. Ries (McDonald Observatory). 0.76-m reflector.

The following orbital elements by B. G. Marsden are from 171 observations, 2003–2008 (mean residual 0′′.7):

Epoch = 2008 May 14.0 TT

T = 2008 May 19.0176 TT  ω = 188.^{o}7935
e = 0.630017  Ω = 66.3890 }2000.0
q = 1.060435 AU  i = 25.5549

a = 2.866176 AU n^{o} = 0.2031187 P = 4.85 years

2008 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green